OpenStudy (anonymous):

What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

wouldnt the answer be 87,384

OpenStudy (anonymous):

thats what i got but im not sure if its right

OpenStudy (callisto):

a=8 r=-2 n=15 sum = a(r^n -1) /(r-1) = 8 [(-2)^15 -1] / [(-2)-1] = 87384 sorry you're right

OpenStudy (anonymous):

alright thanks!

OpenStudy (callisto):

you're welcome